OpenStudy (anonymous):

What is the fourth term in the expansion of (x + 2)^7? Answer 280x^4 120x^3 320x^2 168x^4 What is the fourth term in the expansion of (x + 2)^7? Answer 280x^4 120x^3 320x^2 168x^4 @Mathematics

OpenStudy (cwrw238):

(r+1)th term of (a + b)^n = nCr a^n-r b^r

OpenStudy (anonymous):

look at khayam pascal triangle

OpenStudy (anonymous):

well, it's definitely the one with X^4 :)

OpenStudy (cwrw238):

4th term of (x +2)^7 = 7C3 x^(7-3) 2^3 = 35 x^4 * 8 = 280 x^4